# Unraveling the Connections between Flow Matching and Diffusion Probabilistic Models in Training-free Conditional Generation

### v1

@@ 4 Methodology
Given any ODE in FMs, there exists a probabilistic diffusion path, which could be represented related to the score function in DPMs as follows:
@@
The detailed proof is offered in the supplementary. Proposition 1 indicates one way to explicitly define the score function in the reverse process of the FMs. Thus, we have the following Lemma.
@@
Then, by the Bayes’ rule, we have:
@@
Lemma 1 re-defines a new ODE based on the score function for FMs. In this condition, posterior sampling could work on the FMs again, similar to the posterior sampling on the DPMs.
Posterior sampling for training-free conditional FMs. The next step is to calculate Eq. 11. Under the FMs, $\nabla_{x_{t}}\log p(x_{t})$ and $\nabla_{x_{t}}\log p(c|x_{t})$ could not directly calculate. This is different from the DPMs since we know that $\nabla_{x_{t}}\log p(x_{t})=\epsilon_{\theta}(x_{t},t)$, which is the output of the pre-trained backbone in DPMs. But FMs lose this property.
To achieve posterior sampling, we first implement a math trick for Eq. 11 as follows:
$dx_{t}=[(\lambda x_{t}-\beta_{t}\nabla_{x_{t}}\log p(x_{t}))-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.$ (18)
Then, we could formulate it with a known unconditional term $v(x_{t},t)$ and a conditional term formulation $\nabla_{x_{t}}\log p(c|x_{t})$ by Eq. 10 as follows:
$dx_{t}=[v_{\theta}(x_{t},t)-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.$ (19)
Since $\nabla_{x_{t}}\log p(c|x_{t})$ derives from the probabilistic path, it has the same property and formulation on the DPMs. Following the same estimation from training free conditional DPMs in Eq. 5, we have:
$\nabla_{x_{t}}\log p(c|x_{t})\approx r\nabla_{x_{t}}D(\hat{x}_{0|t},c).$ (20)
Finally, we have to estimate $\hat{x}_{0|t}$ based on the $x_{t}$ via FMs. FMPS offers gradient-aware FMPS (FMPS-gradient) and gradient-free FMPS (FMPS-free) versions by whether $\nabla_{x_{t}}\log p(c|\hat{x}_{0|t})$ needs to introduce the gradient of the diffusion model.
@@
Eq. 21 directly solves the $t\rightarrow 0$ interval to finish the estimation for $x_{t}$. The advantage of gradient-aware estimation is introducing the gradient information of $\frac{\partial v(x_{t},t)}{\partial x_{t}}$ when calculating $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. Gradient information will benefit the conditional tasks [25]. However, it will introduce additional computation costs.
This motivates us to propose our FMPS-free. We re-define the forward process of FMs as follows:
@@
Normalization trick for FMPS. According to the DPMs [25, 7], there is an unstable phenomenon: $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will be over powerful or too weaken compared to the unconditional term, which leading the generated direction to deviate from the data distribution and condition, respectively. To alleviate this, we propose a normalization trick to limit the norm of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. Concretely, ODE decides the strength of $v(\ast,\ast)$ could be measured by its norm since $v(\ast,\ast)$ represents a forward direction in ODE. Thus, we could measure the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. We have:
@@
Eq. 24 limits the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ under the bound of $||v(x_{t},t)||_{2}$, where $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will not too strong and too weaken, thus alleviating the unstable phenomenon.
@@
The overall framework for FMPS based on gradient-aware and gradient-free estimations for $\hat{x}_{0|t}$ are summarized as Algorithm 1 and Algorithm 2, respectively. To sum up, FMPS first leverages the score function in DPMs to re-formulate the ODE of FMs. The score function enables us to introduce the condition directly by changing the marginal distribution to the conditional distribution. In this way, posterior sampling could work for FMs. Thus, we can propose a flow-based posterior sampling to achieve conditional generation.
Additionally, we further introduce a normalization trick to enhance FMPS and improve generation quality. Besides, FMPS also offers a gradient-free estimation for $\hat{x}_{0|t}$, which provides an alternative method to reduce the time costs for calculating posterior sampling.
@@
6:     Calculate $a_{t}$, $b_{t}$, $\dot{a}_{t}$, $\dot{b}_{t}$, $\lambda_{t}$, $v_{\theta}(x_{t},t)$ and $\beta_{t}$. $\triangleright$ Algorithm 1.
@@ 5.1 Implementation Details
Linear inverse problems. Following the previous methods [1], we evaluate FMPS on the CelebA-HQ dataset [26] to test its performance on the linear inverse problems. The linear inverse problems include Inpainting (BOX), Super-resolution (4X), and Gaussian Deblur (kernel size: 61, intensify: 3). The baselines are from the latest methods including Score-SDE [9], ILVR [9], LGD-MC [8], DPS [2], MGPD [3], and DPS+DSG [7]. We use FMPS-free and FMPS-gradient for the linear inverse problem.
Table 3: Ablation study for comparing the FMPS and $\pi$Flow in the linear inverse problem on AFHQ. To make a fair comparison, we choose AFHQ dataset [27] used in $\pi$Flow. We implement $\pi$Flow on the same pre-trained FMs as the FMPS, where the hyperparameters for $\pi$Flow remain in the same setting as reported in its paper. We choose FMPS-gradient since $\pi$Flow relies on the gradient of FMs.
Method SR (4x) Deblur
@@
$\pi$Flow 0.02 0.64 21.22 0.02 0.63 22.02
FMPS-gradient (Our) 0.69 0.13 7.82 0.72 0.12 7.83
Table 4: Ablation study for comparing the FMPS-gradient and FMPS-free based on the SR (4x) on AFHQ.
[…]
FMPS-gradient 0.74 0.12 7.68 15
FMPS-free 0.69 0.13 7.82 7
Non-linear inverse problem. Following the previous methods, we also evaluate FMPS on the CelebA-HQ dataset to test its performance on the non-linear inverse problems. The non-linear inverse problems include Segmenation maps, Sketch, and Face ID. We use the same pre-trained FMs as the linear inverse problems. The baselines are from the latest methods, including Freedom [1], LGD-MC [8], and MPGD [3]. We use FMPS-free and FMPS-gradient for the non-linear inverse problem.
Text-style Generation. Text-style generation guides the generation using text prompts and the given style image. We first evaluate FMPS on the 1000 text-image pairs. The text prompts are randomly selected from PartiPrompts dataset [1]. Then, the style images are from FreeDom, following the previous works. We use FMPS-free for this task to show the potential of FMPS in large FMs since FMPS-gradient cannot afford the GPU cost for calculating the gradient of Stable Diffusion 3.0.
Metrics. All the distance metrics $D(\ast,\ast)$ we use are the same as previous works, where we have listed more details in the supplementary material. For linear inverse problems, we use structural similarity (SSIM) [28], LPIPS [29], and Fréchet inception distance (FID) [30] as the metric to measure its generation quality. For non-linear inverse problems, we use the norm distance between the condition and generated images, FID, and kernel inception distance (KID) [31] as the metrics. Specifically, the distance norm in Face ID tasks is the classifier logit. For the text-style generation, we use the CLIP score [32] and the Gram matrix [1] to measure the distance between images and text prompts and style images, respectively.
@@
Pre-trained Models and settings. To ensure reproduction, we use the pre-trained checkpoint trained in the CelebA-HQ dataset from official Rectified flow [12] for the linear and non-linear inverse problems. For the style task, we use the Stable Diffusion 3 to show the potential of FMPS. The time steps $T$ for all methods in linear and non-linear inverse problems, including FMPS, are set to $T=100$ without specific illustration. All the experiments run on a single RTX 4090 GPU.
Figure 2: The qualitative examples for style tasks using the complex prompts.
@@ 5.2 Qualitative Results
Non-linear inverse problems. We further explore the potential of the FMPS in non-linear inverse problems. We report the comprehensive qualitative evaluation in Table 2. It can be noticed that FMPS-free and FMPS-gradient remain in the high-level generation. Concretely, FMPS-gradient achieves state-of-the-art (SOTA) performance in the segmentation task. Meanwhile, in the sketch task, FMPS-gradient improves by an average 43% improvement in the Distance metric while keeping the lowest FID. Meanwhile, FMPS-free greatly improves the distance metric, although the FID has increased. This proves that the gradient information is useful.
Text-style generation. To prove that the FMPS could work on the real generation tasks, we first report the comprehensive qualitative evaluation in Table 5 using 1000 text-image pairs under nine style types. It can be noticed that FMPS-free could better obey the guide for the text prompts since it achieves the highest CLIP score while keeping the style distance at a low level. Then, to show the ability of FMPS-free, we report the qualitative results using the complex text prompts shown in Fig. 2. It can be noticed that FMPS-free achieves a better generation since it maintains a similar style, following the style images while satisfying the text prompts.
@@ 5.3 Ablation Study
The ablation study first compares the ability between FMPS and $\pi$Flow. Then, we explore the influence of the $\rho$. In the end, we explore the performance of the FMPS-gradient FMPS-free.
Comparison between FMPS and $\pi$Flow. To show the improvement of FMPS compared to $\pi$Flow, we report the comprehensive results shown in Table 3, where the qualitative examples are shown in Fig. 3. It can be seen that the FMPS-gradient achieved SOTA results. Concretely, FMPS improves the FID from 21.22 to 7.82 and 22.02 to 7.83 in SR (4x) and Deblur tasks, respectively. FMPS also significantly improves SSIM and LPIPS metrics. These results prove the flexibility of the FMPS since it could be implemented in pre-trained FMs and maintain high-quality generation.
Figure 3: The qualitative results for comparing FMPS and $\pi$Flow.
Trade-off between FMPS-gradient and FMPS-free. To illustrate the trade-off between FMPS-gradient and FMPS-free, we report the ablation study in Table 4. The advantage of FMPS-free is the low time cost while it achieves 2x speed-up after dropping the gradient. Meanwhile, its performance has decreased at a low level.
@@ 6 Conclusion
We proposed a novel training-free conditional method for FMs called FMPS, where the key step is explicitly defining the score function in the FMs’ ODEs. Thus, the posterior sampling from DPMs could work on the FMs. We further propose FMPS-gradient and FMPS-free to achieve the posterior sampling. The difference is whether we need to calculate the gradient of the FMs during the conditional generation process, where we offer a trade-off solution to reduce the computation cost. The experimental results show that FMPS could better guide the condition while maintaining high-quality generation.
Limitation. There are some constraints for FMPS, similar to the posterior sampling-based works on DPMs. For example, the seed is the key to the generation since if the seed is bad, the process will fail. Then, the color of the generated images will show a slight deviation compared to the DPM methods, which is an interesting finding since SSIM is not well reported in the linear inverse task. This may be due to the influence between stochastic and arbitrary paths. We will work on these further.

### v2

@@ 4.1 The Condition-related Correction Term
Given any ODE in FMs, there exists a valid probabilistic diffusion path, which could be represented related to the score function as follows:
@@
Proposition 1 explicitly defines the score function in the reverse process of the FMs. Thus, we could naturally connect the velocity function of FMs to the score function.
The condition now could be introduced into the velocity function via Bayes’ rule. Thus, we can define our correction term in the following Lemma.
@@
Then, by Bayes’ rule, we have:
@@
Lemma 1 redefines a new ODE based on the score function for FMs. Now, we re-organize Eq 14 as follows:
$\begin{split}dx_{t}&=[(\lambda x_{t}-\beta_{t}\nabla_{x_{t}}\log p(x_{t}))-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt\\ &=[v_{\theta}(x_{t},t)-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.\end{split}$ (21)
It could be found that Eq 21 and Eq. 12 could be matched by $\Delta(x_{t},c)=-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$. Then, $\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ exactly represents the direction from $x_{t}$ to $c$. Thus, we could use $-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ to formulate the correction term by:
$dx_{t}=[v_{\theta}(x_{t},t)-r\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.$ (22)
Eq. 22 proves that the reverse process is always sampled from $p(x|c)$ [12] by $\Delta(x_{t},c)=-r\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$, thereby ensuring the validity of $\Delta(x_{t},c)$.
Compared to Eq. 11, our method is more versatile and can be applied to a wide range of generation tasks, rather than being restricted to linear inversion problems. Moreover, our approach does not require access to all unknown data pairs, making it computationally tractable.
4.2 FMPS
Posterior sampling for training-free conditional FMs. The $\nabla_{x_{t}}\log p(c|x_{t})$ naturally allows the correction to sample from $p(x|c)$, thereby finishing the conditional generation. Our final step is to calculate $\nabla_{x_{t}}\log p(c|x_{t})$, and calculating this term could follow the similar steps of the posterior sampling [17] in SDMs again. Therefore, our framework could be regarded as FMs-based posterior sampling.
Now, we could also approximate $\nabla_{x_{t}}\log p(c|x_{t})$ as [12] by:
$\nabla_{x_{t}}\log p(c|x_{t})\approx\nabla_{x_{t}}D(x_{t},c),$ (23)
where $D(\ast,\ast)$ is the distance metric. Similar to the Freedom [12], posterior sampling enables the use of the energy function (e.g., gradient of the distance metric) to approximate the gradient of the conditional distribution $p(c|x_{t})$. Thus, FMPS enables posterior sampling as flexible as the DPMs-based posterior sampling.
Finally, $x_{t}$ may retain some Gaussian noise to add the bias for the gradient of the distant metric and offer the wrong direction. One way to remove the Gaussian noises is the $x_{0}$ prediction [12] by changing $x_{t}$ to $\hat{x}_{0|t}$ as follows:
$D(x_{t},c)\approx D(\hat{x}_{0|t},c).$ (24)
FMPS offers gradient-aware FMPS (FMPS-gradient) and gradient-free FMPS (FMPS-free) versions to calculate $\hat{x}_{0|t}$ by whether $\nabla_{x_{t}}\log p(c|\hat{x}_{0|t})$ needs to introduce the gradient of $v_{\theta}(x_{t},t)$.
@@ 4.2 FMPS
Eq. 25 directly solves the $t\rightarrow 0$ interval to finish the $x_{t}$ estimation. The advantage of gradient-aware estimation is introducing the gradient information of $\frac{\partial v(x_{t},t)}{\partial x_{t}}$ when calculating $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. Gradient information will benefit the conditional tasks [36]. Thus, it can be regarded as a test-time scaling method [37] since calculating the gradient information introduces extra computational costs.
This motivates us to propose our FMPS-free. We redefine the forward process of FMs as follows:
@@
Normalization trick for FMPS. FM is sensitive to the scale of the direction. FMPS has two directions from $v_{\theta}(x_{t},t)$ and the correction term. We propose a normalization trick to alleviate the scale problem. Concretely, ODE decides the strength of $v(\ast,\ast)$ could be measured by its norm since $v(\ast,\ast)$ represents a forward direction in ODE. Thus, we could measure the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. We have:
@@
Eq. 28 limits the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ under the bound of $||v(x_{t},t)||_{2}$, where $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will not too strong and too weaken.
@@
The overall framework for FMPS based on gradient-aware and gradient-free estimations for $\hat{x}_{0|t}$ is summarized as Algorithm 1 and Algorithm 2, respectively. To sum up, FMPS first proposes the correction term to introduce $c$. Then, FMPS leverages the score function to formulate the correction term, where approximating the correction term returns to the posterior sampling framework again. Therefore, FMPS proposes the two methods to achieve the posterior sampling for FMs.
@@
6:     Calculate $a_{t}$, $b_{t}$, $\dot{a}_{t}$, $\dot{b}_{t}$, $\lambda_{t}$, $v_{\theta}(x_{t},t)$ and $\beta_{t}$.
@@
Figure 2: Quantitative comparison of linear/nonlinear tasks on CelebA-HQ and AFHQ [16]. For linear tasks, we mainly report the results between FMPS and $\Pi$FLow, which is the SOTA method for FMs. For nonlinear tasks, we mainly report the results between FMPS and FreeDom. For FreeDom, we use the optimal setting from UFG [34].
@@ 5.1 Implementation Details
Linear inverse problems. Following the previous methods [12], we evaluate FMPS on the CelebA-HQ dataset [38] to test its performance on the linear inverse problems. The linear inverse problems include Inpainting (BOX), Super-resolution (4X), and Gaussian Deblur (kernel size: 61, intensify: 3). The baselines are from the latest methods, including Score-SDE [7], ILVR [7], LGD-MC [33], DPS [5], MGPD [13], and DPS+DSG [32]. We evaluate FMPS-free and FMPS-gradient on the linear inverse problem.
Table 3: Ablation study for comparing the FMPS-gradient and FMPS-free based on the SR (4x) on AFHQ, where MEM is the memory cost. FMPS-free (w/o) does not use the normalization trick.
@@
FMPS-gradient 0.74 0.12 7.68 15 5.1
FMPS-free 0.69 0.13 7.82 7 3.0
FMPS-free (w/o) 0.55 0.15 41.71 7 3.0
Non-linear inverse problem. Following the previous methods, we also evaluate FMPS on the CelebA-HQ dataset to test its performance on the non-linear inverse problems. The non-linear inverse problems include Segmentation maps, Sketch, and Face ID. We use the same pre-trained FMs as the linear inverse problems. The baselines are from the latest methods, including Freedom [12], LGD-MC [33], and MPGD [13]. We use FMPS-free and FMPS-gradient for the non-linear inverse problem.
Text-style Generation. Text-style generation guides the generation using text prompts and the given style image. We first evaluate FMPS on the 1,000 text-image pairs. The text prompts are randomly selected from PartiPrompts dataset [12]. Then, the style images are from FreeDom, following the previous works. We use FMPS-free for this task to show the potential of FMPS in large FMs since FMPS-gradient cannot afford the GPU cost to calculate the gradient of Stable Diffusion 3.0.
Evaluation Metrics. For linear inverse problems, we use structural similarity (SSIM) [39], LPIPS [40], and Fréchet inception distance (FID) [41] as the evaluation metrics to measure the generation quality. For non-linear inverse problems, we use the norm distance between the condition and generated images, FID, and kernel inception distance (KID) [42] as the metrics. Specifically, the distance norm in Face ID tasks is the classifier logit. For the text-style generation, we use the CLIP score [43] and the Gram matrix [12] to measure the distance between images and text prompts and style images, respectively.
5.2 Distance Metric
All the distance metrics $D(\ast,\ast)$ we use are the same as those in previous works.
5.2.1 Linear Inverse Problems
Inpainting. Given a known center box inpainting mask $m_{i}$ and the Gaussian noise $\epsilon$, the measurement $y$ could be represented as:
$y=m_{i}\odot x+0.05\epsilon,$ (31)
where $x$ is the unknown image. Then, the $d(\ast,\ast)$ is defined as:
$D(\hat{x}_{0|t},y)=||(m_{i}\odot\hat{x}_{0|t}+0.05\epsilon)-y||_{2}.$ (32)
Super-resolution. Given a known interpolation-based super-resolution operator $f_{sr}(\ast)$ (e.g., 4x upsampling), and additive Gaussian noise $\epsilon$, the measurement $y$ can be modeled as:
$y=f_{sr}(x)+0.05\epsilon.$ (33)
Then, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||(f_{sr}(\hat{x}_{0|t})+0.05\epsilon)-y||_{2}.$ (34)
Gaussian Deblur. Given a known Gaussian blur operator $f_{gb}(\ast)$ with 61 kernel size and 3 intensify and the Gaussian noise $\epsilon$, the measurement $y$ could be represented as:
$y=f_{gb}(x)+0.05\epsilon.$ (35)
Then, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||(f_{gb}(\hat{x}_{0|t})+0.05\epsilon)-y||_{2}.$ (36)
5.2.2 Non-linear Inverse Problems
Segmentation. Given a pre-trained face segmentation model $f_{seg}$, the measurement $y$, where the output of $f_{seg}$ is the face segmentation mask map and $y$ is random known segmentation map, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{seg}(\hat{x}_{0|t})-y||_{2}.$ (37)
Sketch. Given a pre-trained image to sketch model $f_{ske}$, the measurement $y$, where the output of $f_{ske}$ is the sketch related to the input image and $y$ is random known sketch, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{ske}(\hat{x}_{0|t})-y||_{2}.$ (38)
Face ID Given a pre-trained face classifier $f_{face}$, the measurement $y$, where the output of $f_{face}$ is the logit value related to the input images and $y$ is random known image, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{face}(\hat{x}_{0|t})-f_{face}(y)||_{2}.$ (39)
5.2.3 Text-Style generation
Given the image encoder from a pre-trained CLIP [43] noted as $E_{i}$ and the style image $y$, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||\text{Gram}(E_{i}(\hat{x}_{0|t}))-\text{Gram}(E_{i}(y))||_{2},$ (40)
where $\text{Gram}(\ast)$ is the gram matrix operator [12].
@@ 5.2.3 Text-Style generation
Pre-trained Models and settings. To ensure reproduction, we use the pre-trained checkpoint trained in the CelebA-HQ dataset from official Rectified flow [1] for the linear and non-linear inverse problems. For the style task, we use Stable Diffusion 3 to show the potential of FMPS. The time steps $T$ for all linear and non-linear inverse problem methods, including FMPS, are set to $T=100$ without specific illustration. All the experiments were run on a single RTX 4090 GPU.
@@ 5.3 Qualitative Results
Non-linear inverse problems. We further explore the potential of the FMPS in non-linear inverse problems. We report the comprehensive qualitative evaluation in Table 2. It can be noticed that FMPS-free and FMPS-gradient remain in the high-level generation. Concretely, FMPS-gradient achieves state-of-the-art (SOTA) performance in the segmentation task. Meanwhile, in the sketch task, FMPS-gradient improves by an average 43% in the Distance metric while keeping the lowest FID. Meanwhile, FMPS-free greatly improves the distance metric, although the FID has increased. This proves that the gradient information is useful.
Text-style generation. To prove that the FMPS could work on the real generation tasks, we first report the comprehensive qualitative evaluation in Table 4 using 1000 text-image pairs under nine style types. It can be observed that FMPS-free more closely adheres to the text prompt guidelines, achieving the highest CLIP score while maintaining a low style distance. Then, to show the ability of FMPS-free, we report the qualitative results using the complex text prompts shown in Fig. LABEL:fig:style. It can be noticed that FMPS-free achieves a better generation since it maintains a similar style, following the style of images while satisfying the text prompts.
Further comparison between FMs and SDMs. To ensure the improvement from FMPS instead of the difference model ability between FMs and SDMs, we make a comparison shown in Table 5. The setting ensure that both methods for FMs and SDMs use the same.
Table 5: Qualitative evaluation of style text-to-image generation. We directly leverage the DPM-based solver, i.e., DPM-solver++ 2M, for FM to make a fair comparison. This solver will first convert the FM to a score-based diffusion paradigm and then implement the DPM-solver. We then implement MPGD in SD3 (MPGD-SD3). Such a setting ensures the same backbone during the training-free condition for both FM and DPMs. The results show that FMPS beats the MPGD and makes a better trade-off. This result also shows the gap that directly implementing the DPMs method on FMs models will degrade the performance of FMs, which enhances the contribution of FMPS.
[…]
MPGD-SD3 439.3 24.34
FMPS-free (Our) 435.2 32.65
5.4 Quantitative Results
Figure 3: Quantitative comparison of complex style text-to-image. We compared all the previous methods on SDMs.
We report the comprehensive quantitative results for both linear/nonlinear (shown in Fig. 2) and style text-to-image (shown in Fig. 3). It could be found that FMPS outperforms the previous methods in all tasks.
@@ 5.5 Ablation Study
We explore the influence of the $r$. Then, we explore the performance of the FMPS-gradient FMPS-free.
Trade-off between FMPS-gradient and FMPS-free. To illustrate the trade-off between FMPS-gradient and FMPS-free, we report the ablation study in Table 3. The advantage of FMPS-free is the low time cost, while it achieves 2x speed-up after dropping the gradient. Meanwhile, its performance has decreased to a low level. In the end, we conducted an ablation study to show the validity of the normalization trick. It can be concluded that the normalization trick is essential for the performance of FMPS. After dropping it, the performance dramatically drops.
@@ 6 Conclusion
We proposed a novel training-free conditional method for FMs called FMPS. FMPS first introduces the condition into the velocity function by introducing a correction term, which offers the additional direction for the condition in the velocity field. Then, we formulate the correction term as the score function in SDMs, enabling the application of posterior sampling within the flow matching framework.
Limitation. There are some constraints for FMPS, similar to the posterior sampling-based works on DPMs. For example, the seed is the key to the generation, since if the seed is bad, the process will fail. Then, the color of the generated images will show a slight deviation compared to the DPM methods, which is an interesting finding since SSIM is not well reported in the linear inverse task. This may be due to the influence of stochastic and arbitrary paths. We will work on these further.